OpenStudy (anonymous):

Solve using the zero factor property: g^2 + 18g + 70 = -11 I got it down to: g^2 + 18g + 59 = 0 - but I can't factor that out.

OpenStudy (asnaseer):

you need to add 11 to both sides. you have subtracted 11 instead.

OpenStudy (anonymous):

Aaaand I feel like an idiot. So, g^2 + 18g + 81 = 0 Let me try to work this out...

OpenStudy (asnaseer):

we ALL make MISKTAKES :) Thats what makes us human.

OpenStudy (anonymous):

I believe it's this...: g^2 + 18g + 81 = 0 (g + 9)(g + 9) = 0 g = -9 Right?

OpenStudy (asnaseer):

correct - well done!

OpenStudy (anonymous):

Thanks! I would've probably never noticed. :D

OpenStudy (asnaseer):

yw